30 November 2007

RHIC Startup:

Yellow tunes were checked at injection, looking for a PS with wrong polarity.  None was found.

BPM offsets were fixed.

H10 magnet controls were modified to allow up to 700 counts change in setting before clipping.  This should allow mode switching without problems.

Stochastic cooling was made operational last night, but Peggy had to modify the tape to get it to work.  The position control bug was fixed (?)  Should be ready to go at next store.  "Go to flattop, push the button, and see if it works" says Mike Blaskowitz.  Beam decay of 20% without cooling turns into 7% with cooling.

Several BLAM dumps pulled the permit over the last few days.  We have been using the conservative approach of including all loss monitors unless they are known bad actors.  Will go to Only known good actors to avoid flakey loss monitors.    Loss monitors should not saturate so that you can make a measurement of whether a dump was clean or dirty.  So may need to add attenuators or reduce gain of some monitors. Should not be using collimtor loss monitors as part of BLAM.

A noisy power supply in Yellow is responsible for the large losses in Yellow at 5:30 this morning.  There will be a two hour access to install diagnostics on this PS in the hope of fixing it this weekend. 

The Air Conditioning in the 5B alcove needs to be turned on.  This could also get done during this morning's access.

BtA does not mode switch correctly.  Currently it needs tweaking(L20) to get from d to Au..

For the weekend, set d intensity to 1E11 and Au intensity to 1E9. Once this is working well at 56 bunches, try increasing the number of bunches.


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020